# Explicit Matching Visual Cues

### v2

@@ 1 Introduction
To bridge this gap, we introduce VLM2-Bench, a benchmark specifically designed to evaluate how well VLMs visually link matching cues. VLM2-Bench is structured around three types of visual cue connection: general cue, person-centric cue, and object-centric cue, encompassing a total of eight subtasks. To balance scalability and quality, we design a semi-automated pipeline with human verification for further refinement. Additionally, our subtasks cover a variety of QA formats—including T/F, multi-choice, numerical, and open-ended questions—totaling over 3,000 question-answer pairs. To better evaluate model performance, we also design specific metrics tailored to various task.
We conduct a comprehensive evaluation of 8 open-source models and GPT-4o on our VLM2-Bench. Despite VLMs generally possessing extensive knowledge, some models perform on par with, or even worse than, the chance-level baseline on our vision-centric tasks. Notably, GPT-4o also underperforms, lagging behind human-level accuracy by 34.80%. This highlights the significant room for improvement in VLMs’ ability to link visual cues. Furthermore, we introduce various language-side and vision-side prompting techniques to explore whether they can enhance the models’ performance on the benchmark. Through experimental results and case studies, we present eight key observations, hoping that these insights will guide future improvements in VLMs for vision-centric tasks.
@@ 2 VLM2-Bench
VLM2-Bench is a benchmark designed to assess models’ ability to visually link matching cues when processing multiple images or videos. This section introduces the three main categories of VLM2-Bench—general cue2.1), object-centric cue2.2), and person-centric cue2.3)—detailing their associated subtasks, data collection process, and QA pair construction.
@@ 3.1 Metric Design
$Acc_{num}=1-\frac{1}{n}\sum_{i=1}^{n}w_{i}\cdot\epsilon_{i}^{-\alpha}.$ (4)
@@ Methods.
(a) Results of CoT-normal, CoT-special, and VP-grid on GC.
(b) Results of CoT and VP-zoom-o on OC.
(c) Results of CoT and VP-zoom-p on PC.
@@ 5 Related Work
(Hurst et al., 2024; Team, 2025; Zhang et al., 2024a; Li et al., 2024b; Ye et al., 2024a; Chen et al., 2024b; Liang et al., 2024b) have significantly broadened their capabilities. Previously restricted to processing single-image inputs, many VLMs can now handle multi-image and even video inputs, allowing them to capture richer and more dynamic visual contexts. Additionally, with access to a growing volume of high-quality visual-textual paired training data (Pi et al., 2024b; Garg et al., 2024; Chen et al., 2023; Zhang et al., 2024c; Wang et al., 2024c), these models have shown substantial improvements in perceiving subtle visual cues and their relationships, enabling them to engage in more nuanced reasoning about visual content. Furthermore, VLMs are increasingly applied in real-world scenarios, including navigation (Weerakoon et al., 2024), planning (Yang et al., 2024), and autonomous driving (Jiang et al., 2024), solidifying their role in bridging vision and language for practical applications. However, to truly integrate into everyday life, VLMs still have significant room for improvement when it comes to more fundamental but common visual tasks, such as those assessed in our benchmark.
@@
plays a critical role in guiding their future development (Liang et al., 2024a; Yin et al., 2023; Chen et al., 2024a). These benchmarks typically focus on assessing the models’ fine-grained perception (Li et al., 2024a; Tong et al., 2024), reasoning abilities (Lu et al., 2022; Yu et al., 2023; Wu et al., 2024), commonsense knowledge (Yue et al., 2024). In addition, evaluations targeting multi-image and video inputs are designed to measure the new competencies that VLMs require as their visual context extends. These tasks include captioning (Yue et al., 2024; Yu et al., 2019), retrieval (Wang et al., 2024a; Li et al., 2025), comparison (Wu et al., 2025; Jiao et al., 2024), and temporal reasoning (Liu et al., 2024b). However, existing benchmarks focus on evaluating VLMs’ ability to interpret visual cues based on their knowledge. In contrast, humans typically solve such tasks by explicitly matching visual cues without relying on extensive background knowledge. To better assess whether they can replicate this human-like ability, we propose VLM2-Bench, which focuses on linking and matching explicit visual cues.
@@ 7 Conclusion
VLM2-Bench focuses on evaluating visual cue linking but does not cover all possible scenarios. Additionally, while it provides valuable insights, its scale is limited, and model performance may not fully generalize to all real-world settings. Automated evaluation constraints limit the inclusion of open-ended questions in our benchmark, impacting the assessment of models’ vision-centric reasoning abilities. Expanding task diversity and refining evaluation methods remain important directions for future work.
@@ D.1 Chance-level
$E(L)=1-\frac{1}{L^{2}}\sum_{N=1}^{L}\sum_{\hat{N}=1}^{L}w(L)\cdot\epsilon(N,\hat{N})^{-\alpha},$
@@ E.2 OC (Object-centric Cue)
We predefine 8 types of common objects, with each type containing 5 meta-objects. For each meta-object, we collect four images that represent the same object from different angles and scene conditions.
@@ F.1 Prompts for LLM-as-Evaluator
When models answer our free-form PC-VID questions, their responses are evaluated by GPT-4o using the scoring prompts detailed in Tables 17 and 18. Specifically, for videos following a $\mathcal{P}\rightarrow\neg\mathcal{P}$ sequence, GPT-4o assesses whether the model explicitly distinguishes that the first individual ($\mathcal{P}$) and the second individual ($\neg\mathcal{P}$) are different. In this case, if the model successfully makes this distinction, it receives a score of 1; otherwise, it is given a score of 0.
@@ G.2 Case for VP-grid in General Cue Task
A example failure case in VP-grid.
@@
A example of success case in VP-grid.

### v3

@@ 1 Introduction
To bridge this gap, we introduce VLM2-Bench, a benchmark specifically designed to evaluate how well VLMs visually link matching cues. VLM2-Bench is structured around three types of visual cue connection: general cue, person-centric cue, and object-centric cue, encompassing a total of eight subtasks. To balance scalability and quality, we design a semi-automated pipeline with human verification for further refinement. Additionally, our subtasks cover a variety of QA formats—including T/F, multi-choice, numerical, and open-ended questions—totaling over 3,000 question-answer pairs. To better evaluate model performance, we also design specific metrics tailored to various tasks.
We conduct a comprehensive evaluation of 9 open-source models and GPT-4o on our VLM2-Bench. Despite VLMs generally possessing extensive knowledge, some models perform on par with, or even worse than, the chance-level baseline on our vision-centric tasks. Notably, GPT-4o also underperforms, lagging behind human-level accuracy by 34.80%. This highlights the significant room for improvement in VLMs’ ability to link visual cues. Furthermore, we introduce various language-side and vision-side prompting techniques to explore whether they can enhance the models’ performance on the benchmark. Through experimental results and case studies, we present eight key observations, hoping that these insights will guide future improvements in VLMs for vision-centric tasks.
@@ 2 VLM2-Bench
As shown in Figure 2, VLM2-Bench is a benchmark designed to assess models’ ability to visually link matching cues when processing multiple images or videos. This section introduces the three main categories of VLM2-Bench—general cue2.1), object-centric cue2.2), and person-centric cue2.3)—detailing their associated subtasks, data collection process, and QA pair construction.
@@ 3.1 Metric Design
$Acc_{num}=1-\frac{1}{n}\sum_{i=1}^{n}w_{i}\cdot\epsilon_{i}^{\alpha}.$ (4)
@@ Methods.
(a) CoT-normal, CoT-special, and VP-grid on GC.
(b) CoT and VP-zoom-o on OC.
(c) CoT and VP-zoom-p on PC.
@@ 5 Related Work
(Hurst et al., 2024; Team, 2025; Zhang et al., 2024a; Li et al., 2024b; Ye et al., 2024a; Chen et al., 2024b; Liang et al., 2024b) have significantly broadened their capabilities. Previously restricted to processing single-image inputs, many VLMs can now handle multi-image and even video inputs, allowing them to capture richer and more dynamic visual contexts. Additionally, with access to a growing volume of high-quality visual-textual paired training data (Pi et al., 2024b; Garg et al., 2024; Chen et al., 2023; Zhang et al., 2024c; Wang et al., 2024c), these models have shown substantial improvements in perceiving subtle visual cues and their relationships, enabling them to engage in more nuanced reasoning about visual content. Furthermore, VLMs are increasingly applied in real-world scenarios, including navigation (Weerakoon et al., 2024), planning (Yang et al., 2024), autonomous driving (Jiang et al., 2024), and report generation Reddy et al. (2024), solidifying their role in bridging vision and language for practical applications. However, to truly integrate into everyday life, VLMs still have significant room for improvement when it comes to more fundamental but common visual tasks, such as those assessed in our benchmark.
@@
plays a critical role in guiding their future development (Liang et al., 2024a; Yin et al., 2023; Chen et al., 2024a). These benchmarks typically focus on assessing the models’ fine-grained perception (Li et al., 2024a; Tong et al., 2024), reasoning abilities (Lu et al., 2022; Yu et al., 2023; Huang et al., 2025), and commonsense knowledge (Yue et al., 2024; Wu et al., 2024). In addition, evaluations targeting multi-image and video inputs are designed to measure the new competencies that VLMs require as their visual context extends. These tasks include captioning (Yue et al., 2024; Yu et al., 2019), retrieval (Wang et al., 2024a; Li et al., 2025b), comparison (Wu et al., 2025; Jiao et al., 2024), and temporal reasoning (Liu et al., 2024b). However, existing benchmarks focus on evaluating VLMs’ ability to interpret visual cues based on their knowledge. In contrast, humans typically solve such tasks by explicitly matching visual cues without relying on extensive background knowledge. To better assess whether they can replicate this human-like ability, we propose VLM2-Bench, which focuses on linking and matching explicit visual cues.
@@ 7 Conclusion
VLM2-Bench focuses on evaluating visual cue linking but does not cover all possible scenarios. Additionally, while it provides valuable insights, its scale is limited, and model performance may not fully generalize to all real-world settings. Automated evaluation constraints limit the inclusion of open-ended questions in our benchmark, impacting the assessment of models’ vision-centric reasoning abilities. Expanding task diversity and refining evaluation methods remain important directions for future work. In future research, model self-play Li et al. (2025a), self-correction He et al. (2024), and synthetic data pretraining Qin et al. (2025) may also be interesting to explore.
@@ D.1 Chance-level
$E(L)=1-\frac{1}{L^{2}}\sum_{N=1}^{L}\sum_{\hat{N}=1}^{L}w(L)\cdot\epsilon(N,\hat{N})^{\alpha},$
@@ E.2 OC (Object-centric Cue)
We predefine 8 types of common objects, with each type containing 5 meta-objects to ensure a class-balanced sampling and avoid long-tail distribution (Yao and Li, 2023). For each meta-object, we collect four images that represent the same object from different angles and scene conditions.
@@ F.1 Prompts for LLM-as-Evaluator
When models answer our free-form PC-VID questions, their responses are evaluated by an evaluator model (Yao et al., 2024) (here GPT-4o) using the scoring prompts detailed in Tables 17 and 18. Specifically, for videos following a $\mathcal{P}\rightarrow\neg\mathcal{P}$ sequence, GPT-4o assesses whether the model explicitly distinguishes that the first individual ($\mathcal{P}$) and the second individual ($\neg\mathcal{P}$) are different. In this case, if the model successfully makes this distinction, it receives a score of 1; otherwise, it is given a score of 0.
@@ G.2 Case for VP-grid in General Cue Task
An example failure case in VP-grid.
@@
An example of success case in VP-grid.